OpenStudy (anonymous):

1)What are the first five terms of the sequence given by the formula an=4n+1 ? a. 1,5 ,9,13,17 b. 9,13,17,21,25 c. 5,9,13,17,21 d. 4,8,12,16,20 2)Write a recursive definition for the sequence 8,6,4,2 ... a. a1=8, an=an-1-2 b. a1=2, an=an-1+6 c. a1=8,an=an-1+2 d. a1=2, an=an-1+8

OpenStudy (anonymous):

\[ a_1= 4(1)+1=5\\ a_2=4(2) +1=9\\ a_3= 4(3) +1=13 \] Which one is it?

OpenStudy (anonymous):

i will let you know as soon as I submit my answers! Thank you!

OpenStudy (anonymous):

Can't you find out without submitting your answer? Which sequence in your choices start with 5,9,13

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

No A starts with 1, 5, 9, 13

OpenStudy (anonymous):

yes, in the work you gave me you put 1,5,9,13

OpenStudy (anonymous):

Read my post again \[ a_1= 4(1)+1=5\\ a_2=4(2) +1=9\\ a_3= 4(3) +1=13 \]

OpenStudy (anonymous):

It is 5,9,13